OpenStudy (kakarot):

solve using subtitution y = 2x – 8 y = -x + 1

OpenStudy (onepieceftw):

Substitute one of your equations for a variable in the other. To make it simple substitute 2x-8 for y in y=-x+1

OpenStudy (onepieceftw):

\[(2x - 8) = -x + 1\]

OpenStudy (kakarot):

ok

OpenStudy (onepieceftw):

subtract x from both sides then we have x - 8 = 1

OpenStudy (onepieceftw):

add 8 to both sides to get x = 9

OpenStudy (onepieceftw):

Now put nine into either of the original equations to find y

OpenStudy (kakarot):

3, -2?

OpenStudy (onepieceftw):

y = 2(9) - 8 y = 10